# Quantum reservoir computing for predicting and characterizing chaotic maps

### v1

@@ Abstract
Robustness of forecasting.– So far, our results have assumed perfect unitary evolution, as well as randomized but fixed instances of reservoir Hamiltonian initialization. However, in the NISQ era, the former is a simplistic assumption as the presence of noise leads to non-unitary evolution in real setups. Moreover, the appeal of reservoir computing lies in the fact that the algorithm should work well regardless of the exact internal coupling parameters of the reservoir. Thus, for actual implementation, we must account for both these factors to benchmark the realistic predictive accuracy of our protocol. This analysis has been performed in the end matter below. Firstly, we show that by training the system with the measured data from a noisy reservoir we can reach significant robustness against dephasing. Secondly, our statistical analysis demonstrates that our protocol is robust against randomness in reservoir couplings.
Conclusion.– In this work, we have proposed a QRC algorithm for predicting and characterizing nonlinear dynamics in discrete maps with chaotic regimes. We have demonstrated the effectiveness of our method on the logistic and Hénon maps. While we can achieve accurate prediction about the future steps of our time series, we have also been able to optimize the QRC encoding protocol based on the features of nonlinear maps studied herein, paving the way for further optimization of the QRC architecure for experimental time series data. In addition, the proposed QRC algorithm is resilient against decoherence and reservoir randomness, making it suitable for near-term quantum devices. We note here that in contrast to other recent proposals of applying QRC to chaotic dynamics forecasting [52, 53] studying continuous dynamical evolution, we study discrete chaotic maps. Moreover, we use simple phase encoding as opposed to computationally costly amplitude encoding [52]; and employ a simple reservoir only requiring nearest neighbor XX interaction which is readily available in superconducting quantum simulators [54, 55, 56, 57], ion-trap systems [58, 59, 60] and optical lattices [61, 62, 30] as opposed to requiring complex reservoir engineering with fully-connected [53] or non-sparse Hamiltonians [52].
Acknowledgments.– AB acknowledges support from the National Natural Science Foundation of China (grants No. 12274059, No. 12574528 and No. 1251101297). LM gratefully acknowledges the support of the “Hundred Talents” program of the University of Electronic Science and Technology of China, the “Outstanding Young Talents Program (Overseas)” of the National Natural Science Foundation of China, and the talent programs of the Sichuan province and Chengdu municipality.
I End Matter
In this end matter, we discuss the robustness of the QRC protocol.
I.1 Robustness against decoherence
@@
I.2 Robustness against reservoir randomness
@@
In this section, we discuss the robustness of our QRC protocol for predicting chaotic maps with respect to the choice of the parameters in the quantum reservoir. In general, the lack of sensitivity to such parameterization is a highly desirable trait for reservoir computing, where the explicit goal is to only adjust the weight matrix W at the end. To check this, we randomly selected $1000$ Hamiltonians for both the logistic and the Hénon maps and binning their respective predictive errors, represented by RMSEs, in the histograms of Fig. 5. The results indicate that the majority of Hamiltonians yield low RMSE, with only a small fraction exhibiting big errors. Moreover, in both cases, the histograms are Poissonian in nature, which means only a very small fraction of all the Hamiltonians are ‘problematic’ with the vast majority of randomly chosen Hamiltonians yielding close to maximum predictive accuracy. This is consistent with the measure concentration phenomenon encountered in locally entangled many-body systems such as the present reservoir based on disordered transverse XY spin chains, where most dynamical states are in fact concentrated within a small subspace of the overall Hilbert space  [64, 65, 66]. While an explicit quantification is beyond the scope of the present work, it appears randomly sampled Hamiltonians do typically arise from this subspace and the exceptions outside this subspace are rare, thus a Poissonian treatment is apt.

### v2

@@ VI.1 Open-loop prediction of QRC
VI.2 Closed-loop prediction of QRC
In the previous section, we found that performing multi-step-ahead direct prediction can be demanding in terms of quantum resources. For instance, one typically needs a larger number of input repetitions $n_{\mathrm{rep}}$ to induce higher-order nonlinear features, which becomes crucial when forecasting further into the future. In this section, we investigate how to combine QRC with closed-loop forecasting to progressively predict future values of the logistic and Hénon maps. In a closed-loop setting, the model’s own predictions are fed back as inputs for subsequent steps, thereby generating a multi-step trajectory through iterative one-step-ahead prediction. Specifically, consider a QRC model with $d$ layers. Given an initial input window $(x_{t-d},x_{t-d+1},\ldots,x_{t-1})$, our goal is to predict the value at time t1+kt{-}1{+}k , with $k$ being the prediction horizon. Rather than training the QRC to output $x_{t-1+k}$ directly (as in the direct multi-step setup for k=2k{=}2 discussed earlier), we first use the model to produce a one-step prediction $\hat{x}_{t}$. We then slide the input window forward to $(x_{t-d+1},\ldots,x_{t-1},\hat{x}_{t})$, and use the same model to output the next prediction $\hat{x}_{t+1}$. This procedure is repeated, updating the input window at each step, until the model produces $\hat{x}_{t-1+k}$. This closed-loop configuration is closer to realistic forecasting scenarios. However, because it relies on feeding back predicted values, the task typically becomes more challenging as the prediction horizon $k$ increases. This can be attributed to successive magnification of prediction errors at each step arising out of use of previously simulated predictions as training data for further prediction. Since closed-loop forecasting is implemented via iterative one-step prediction, we adopt the optimal QRC model that yields the best single-step performance. For the logistic map, our previous results show that the optimal configuration is $d=2$ and $n_{\mathrm{rep}}=2$ (see Fig. 3(a)). For the Hénon map, our previous results show that the optimal configuration is $d=1$ and $n_{\mathrm{rep}}=2$ (see Fig. 3(b). Therefore, we use these settings respectively for logistic and Hénon maps throughout, aiming to minimize the error introduced at each step and thereby maintain reliable long-horizon forecasts.
Figure 4: Actual (red) vs predicted (blue) bifurcation diagram of the logistic map with different closed-loop prediction horizons $k$, where for (a) $k=3$, (b) $k=5$, (c) $k=7$ and (d) $k=9$. Actual (red) vs predicted (blue) bifurcation diagram of the Hénon map with different closed-loop prediction horizons $k$, where for (e) $k=3$, (f) $k=5$, (g) $k=7$, and (h) $k=9$.
In Fig. 4, we plot the bifurcation diagram of the logistic map (Figs. 4(a)-(d)) and the Hénon map (Figs. 4(e)-(h)) obtained from multi-step forecasting using the closed-loop strategy. We display the predicted asymptotic values $x_{\infty}$ by again collecting the points from $x_{151}$ to $x_{200}$. We consider four different prediction horizons, namely k=3,5,7,9k{=}3,5,7,9. From Fig. 4, we observe that the model can accurately reproduce the broad bifurcation diagram of the logistic map with respect to the control parameter $r$ (logistic map), $a$ (Hénon map) in the fixed-point regime as well as across the periodic windows. For the logistic map, for horizon k=3k{=}3, the QRC prediction yields an excellent reconstruction of the bifurcation structure even in the chaotic regime. However, when the prediction horizon increases to k=5k{=}5 and beyond, the performance deteriorates in the strongly chaotic region close to $r\simeq 4$. In this regime, the iterative feedback inherent to closed-loop forecasting leads to gradual accumulation of errors, and the predicted values deviate substantially from the valid range of the logistic map, resulting in poor agreement with the true bifurcation pattern. This behavior reflects a fundamental challenge of long-horizon prediction in chaotic systems, where small inaccuracies are exponentially amplified. For the Hénon map, the same pattern is seen, although here the disagreement between predicted values and the ground-truth is even more pronounced, which we may ultimately ascribe to the fact that the Hénon map is a two-dimensional map with errors accumulating along both variables. At the same time, it is worth emphasizing that our QRC employs only d=2d{=}2 layers for the logistic map and only d=1d{=}1 layer for the Hénon map. Consequently, even for k=3{k=}3, once the prediction is rolled forward, the model necessarily relies entirely on its own intermediate predictions (e.g., $\hat{x}_{t}$ and $\hat{x}_{t+1}$) without access to any further ground-truth values. This highlights both the capability as well as the limitation of our QRC to perform genuine multi-step forecasting within a closed-loop setting.
For a quantitative understanding of the prediction error for the closed-loop strategy, we once again employ the RMSE of prediction error as a function of the map parameter $r$ (logistic map) and $a$ (Hénon map) for various prediction horizons $k$. In Fig.  5(a) ((c)), we plot the largest Lyapunov exponent (LLE) together with the prediction error (RMSE) as functions of the control parameter rr($a$) of the logistic map (Hénon map) for different closed-loop prediction horizons $k$. We observe that before the LLE first becomes positive, i.e., before the logistic map enters the chaotic regime, the QRC achieves very similar performance across different prediction horizons $k$, all with uniformly small prediction errors. However, as $r$ increases further and the map enters the chaotic regime via successive bifurcations, the prediction error gradually grows for all prediction horizons $k$. Nevertheless, when $r$ lies within islands of stability (periodic windows embedded in chaos), the RMSE exhibits a pronounced dip compared to the surrounding chaotic region. This indicates that the QRC is highly sensitive to such stable windows and can effectively distinguish them from fully chaotic dynamics. Overall, the trend of prediction error versus logistic map parameter $r$ in the closed-loop strategy is highly consistent with the single-step prediction results reported earlier (see Fig. 2). At the same time, Fig. 5(a) also indicates the prediction error increases with the forecasting horizon $k$, which can be attributed to the accumulation of prediction errors inherent in the closed-loop iterative procedure. All these results described above for the logistic map indicating progressively worsening prediction accuracy with onset of chaos qualitatively agree with the results obtained for the Hénon map in Fig. 5(c).
Figure 5: (a) RMSE of predicting bifurcation diagram of the logistic map with different horizon steps $k=3,5,7,9$. LLE (black line) and chaotic regimes (grey) are depicted for reference. (b) RMSE of prediction at different regimes $r=2.9$ (fixed point, blue), $r=3.5$ (4-periodic orbit, teal), $r=3.829$ (island of stability, magenta), $r=4.0$ (chaotic, red) with prediction horizon $k$ ranging from 1 to 10. (c) same as (a) but for Henon map. (d) same as (b) but for Henon map at different regimes $a=1.005$ (4-periodic orbit, teal) , $a=1.25$ (7-periodic orbit, blue), $a=1.4$ (chaotic, magenta).
To further examine the closed-loop performance of the QRC across different prediction horizons, we select four representative values of the control parameter $r$, namely: (i) r=2.9r{=}2.9, where the logistic map converges to a single fixed point; (ii) r=3.5r{=}3.5, where the logistic map orbit exhibits periodic oscillations between four values; (iii) $r=3.829$, corresponding to an island of stability; and (iv) $r=4$, where the map is fully chaotic. For each of these cases, we generate closed-loop predictions for the time interval $x_{151}$ to $x_{200}$ under different horizons, and we repeat the test over $10$ different initial conditions. The results are reported in Fig. 5(b), where both the prediction-error and its variance (error-bars) have been depicted with respect to various prediction horizons $k$ ranging from k=1k{=1} to k=10k{=}10. For the single fixed point $r=2.9$, the QRC achieves consistently accurate forecasts across all horizons, with errors $\sim 10^{-4}$. For the 4-periodic oscillation regime at $r=3.5$, where the orbit oscillates among four discrete values, the QRC likewise maintains high accuracy for all prediction horizons considered with errors $\sim 10^{-3}$. At $r=3.829$ (island of stability), the prediction error increases moderately with the prediction horizon $k$, typically from $\sim 10^{-2}~(k{=}1)$ towards $\sim 10^{-1}~(k{=}10)$. In contrast, for the fully chaotic case $r=4$, the forecasting error grows significantly as the horizon increases, rising from $\sim 10^{-2}~(k{=}1)$ up to $\sim 10^{0}~(k{=}6)$. For completeness, in appendix A together with Fig. 9 therein, we provide explicit comparisons between the QRC predictions and the corresponding ground-truth trajectories for prediction horizons k=3,5,7,9k{=}3,5,7,9 at each of the four representative values of $r$ listed above. The corresponding results for the Hénon map are depicted in Fig. 5(d) for various values of the map parameter $a$ - namely, 4-periodic orbits ($a=1.005$), 7-periodic orbits ($a=1.25$), and chaotic regime ($a=1.4$). These results again agree qualitatively with the results for the logistic map, viz., prediction error increases even more significantly with the prediction horizon $k$ within the chaotic regime. This once again highlights the intrinsic difficulty of long-horizon prediction in chaotic systems under closed-loop strategy.
From the results above, we conclude that using the closed-loop forecasting strategy, our QRC model can capture the distinctive properties of the logistic map across a wide range of dynamical regimes. In particular, when the logistic map is in non-chaotic regimes, the model can accurately predict the values of $x_{t}$ even up to relatively large prediction horizons, indicating its ability to capture the long-term behavior of the map in regimes where trajectories are stable or periodic. In the chaotic regime, the model can still reproduce the dynamics reasonably well for short prediction horizons. However, as the prediction horizon $k$ increases, accurate numerical prediction becomes extremely challenging due to the intrinsic sensitivity of chaotic dynamics and the accumulation of errors in the closed-loop strategy. Overall, our results demonstrate that the proposed QRC model possesses long-horizon closed-loop forecasting capability despite using a small input window and very low circuit depth (only two layers). In addition, the behavior of error in long-horizon closed-loop prediction can be used for discriminating between chaotic and non-chaotic behaviors.
VII Robustness of forecasting
@@
So far, our results have assumed perfect unitary evolution, as well as randomized but fixed instances of reservoir Hamiltonian initialization. However, in the NISQ era, the former is a simplistic assumption as the presence of noise leads to non-unitary evolution in real setups. Moreover, the appeal of reservoir computing lies in the fact that the algorithm should work well regardless of the exact internal coupling parameters of the reservoir. Thus, for actual implementation, we must account for both these factors to benchmark the realistic predictive accuracy of our protocol. Moreover, the final expectation values readout theoretically requires infinite trials on identical ensembles to fully reconstruct corresponding probabilities, while practically we are firmly in the finite-sampling regime. Thus, firstly, we show that by training the system with the measured data from a noisy reservoir we can reach significant robustness against dephasing. Secondly, we demonstrate with statistical analysis that our protocol is robust against randomness in reservoir couplings. Finally, we show that the a relatively small number of measurements for readout already suffices to agree with theoretical results reported here.
VII.1 Robustness against decoherence
@@
Figure 8: Norm $W^{*}||$ of optimal weight matrix $W^{*}$ of QRC for the logistic map (left), and the Hénon map (right).
VII.2 Robustness against reservoir randomness
@@
In this section, we discuss the robustness of our QRC protocol for predicting chaotic maps with respect to the choice of the parameters in the quantum reservoir. In general, the lack of sensitivity to such parameterization is a highly desirable trait for reservoir computing, where the explicit goal is to only adjust the weight matrix W at the end. To check this, we randomly selected $1000$ Hamiltonians for both the logistic and the Hénon maps and binning their respective predictive errors, represented by RMSEs, in the histograms of Fig. 7. The results indicate that the majority of Hamiltonians yield low RMSE, with only a small fraction exhibiting big errors. Moreover, in both cases, the histograms are Poissonian in nature, which means only a very small fraction of all the Hamiltonians are ‘problematic’ with the vast majority of randomly chosen Hamiltonians yielding close to maximum predictive accuracy. This is consistent with the measure concentration phenomenon encountered in locally entangled many-body systems such as the present reservoir based on transverse XY spin chains, where most dynamical states are in fact concentrated within a small subspace of the overall Hilbert space  [59, 60, 61]. While an explicit quantification is beyond the scope of the present work, it appears randomly sampled Hamiltonians do typically arise from this subspace and the exceptions outside this subspace are rare, thus a Poissonian treatment is apt.
VII.3 Robustness against finite sampling error
In the numerical simulations presented above, we directly computed the expectation values of Pauli-$X$ observables at the last layer. However, these expectation values are theoretically only valid in the asymptotic limit of large number of trials. In an experimental setting, however, these expectation values must be estimated from a finite number of measurement shots. Assuming that each observable $X$ is measured with $M$ shots and independent measurements on identically prepared samples, the estimated expectation value $\langle X\rangle_{\mathrm{sam}}$ is centered around the true expectation value $\langle X\rangle$ with an width $\Delta(X)$, i.e.,
$\langle X\rangle=\langle X\rangle_{\mathrm{sam}}\pm\Delta(X),$ (6)
where the statistical estimation error $\Delta(X){\sim}\sqrt{\frac{\langle X^{2}\rangle-\langle X\rangle^{2}}{M}}$ diminishes as ${\sim}\mathcal{O}(1/\sqrt{M})$. We note that since the Pauli-$X$ operator satisfies $\langle X^{2}\rangle=1$ and $0\leq\langle X\rangle^{2}\leq 1$ throughout, the numerator $\langle X^{2}\rangle-\langle X\rangle^{2}$ is bounded everywhere.
We now note that the prediction from the QRC is obtained via a linear readout of the quantum measured features, i.e.,
$\hat{y}=W\vec{m},$ (7)
where $\vec{m}{=}[\langle X_{1}\rangle,\langle X_{2}\rangle,\cdots,\langle X_{n_{\textrm{I}}+n_{\textrm{H}}}\rangle]^{T}$. Writing $\vec{m}=\vec{m}_{\mathrm{sam}}\pm\vec{m_{\Delta}}$, where $\vec{m_{\Delta}}{=}[\Delta(X_{1}),\Delta(X_{2}),\cdots,\Delta(X_{n_{\textrm{I}}+n_{\textrm{H}}})]^{T}$ yields
$\hat{y}=W(\vec{m}_{\mathrm{sam}}\pm\vec{m_{\Delta}}).$ (8)
In other words, the finite-sampling error in the measured feature vector $\vec{m}$ propagates linearly into the prediction. Practically, we want the prediction deviation induced by this finite sampling error to be below some tolerance $\epsilon$, namely:
$\|W\vec{m}_{\mathrm{sam}}-W\vec{m}\|_{2}=\|W\vec{m_{\Delta}}\|_{2}<\epsilon.$ (9)
Now, vide the Cauchy-Schwarz inequality, we already know that
$\|W\vec{m_{\Delta}}\|_{2}\leq\|W\|_{2}\|\vec{m_{\Delta}}\|_{2}<\epsilon.$ (10)
Because $\|\vec{m_{\Delta}}\|_{2}=\sqrt{\sum_{i}^{n_{\textrm{I}}+n_{\textrm{H}}}\frac{1-\langle X_{i}\rangle}{M}}\leq\sqrt{\frac{n_{\textrm{I}}+n_{\textrm{H}}}{M}}$, this inequality yields an explicit relationship between the required shot number $M$ and the desired tolerance $\epsilon$ as $\frac{(n_{\textrm{I}}+n_{\textrm{H}})\|W\|_{2}}{\epsilon^{2}}\leq M$. In particular, this analysis indicates that it is advantageous for the readout weights to have a moderate norm; i.e., the norm $\|W^{*}\|^{2}$ corresponding to the optimal weight matrix $W^{*}$ arrived at from the training should not be too large in order to avoid excessive amplification of finite-sampling error. In Fig. 8, we plot $\|W^{*}\|^{2}$ as a function of map parameters ($r$ for the logistic map and $a$ for the Hénon map). In both cases, We find that $\|W^{*}\|^{2}$ remains modest across the parameter range, suggesting that an experimentally feasible number of measurement shots should be sufficient to overcome the finite sampling error.
Thus, we have demonstrated the robustness of QRC against coherent and incoherent noise, as well as finite sampling errors. At the same time, we note that several recent works on QRC have aimed to further enhance noise resilience by incorporating error-mitigation-inspired strategies, for example, by modifying the training protocol [62] and/or the measurement scheme [43]. This line of research is complementary to the results presented here, as such techniques may further improve the predictive accuracy and stability of QRC, especially for long-horizon forecasting tasks.
VIII Conclusion
In this work, we have proposed a QRC algorithm for predicting and characterizing nonlinear dynamics in discrete maps with chaotic regimes. We have demonstrated the effectiveness of our method on the logistic and Hénon maps. While we can achieve accurate prediction about the future steps of our time series, we have also been able to optimize the QRC encoding protocol based on the features of nonlinear maps studied herein, paving the way for further optimization of the QRC architecure for experimental time series data. In addition, through closed-loop prediction of distant future steps one can discriminate non-chaotic versus chaotic phases. In this situation, while the closed-loop strategy can faithfully predict distant future outcomes in non-chaotic phases, in the chaotic phase the error grows significantly as the horizon for future steps increases. The proposed QRC algorithm is resilient against decoherence and reservoir randomness, making it suitable for near-term quantum devices, and relies on reasonably small number of measurement samples. We note here that in contrast to other recent proposals of applying QRC to chaotic dynamics forecasting [63, 64] studying continuous dynamical evolution, we study discrete chaotic maps. Moreover, we use simple phase encoding as opposed to computationally costly amplitude encoding [63]; and employ a simple reservoir only requiring nearest neighbor XX interaction which is readily available in superconducting quantum simulators [65, 66, 67, 68], ion-trap systems [69, 70, 71] and optical lattices [72, 73, 33] as opposed to requiring complex reservoir engineering with fully-connected [64] or non-sparse Hamiltonians [63].
IX Acknowledgments
AB acknowledges support from the National Natural Science Foundation of China (grants No. 12274059, No. 12574528, No. 1251101297 and No. W2541020). LM gratefully acknowledges the support of the “Hundred Talents” program of the University of Electronic Science and Technology of China, the “Outstanding Young Talents Program (Overseas)” of the National Natural Science Foundation of China, and the talent programs of the Sichuan province and Chengdu municipality.
Appendix A Closed-loop prediction with different prediction horizons for logistic and Hénon maps
In this Appendix, we report the prediction results for the logistic and Hénon maps under different closed-loop horizons (see Fig. 9). For the logistic map, we focus on four representative values of the control parameter $r$ ($r=2.9,3.5,3.829,4$ as specified in the main text). For the Hénon map, we focus on three representative values of the control parameter $a$ ($a=1.005,1.25,1.4$ as specified in the main text). For the non-chaotic cases of both maps, the QRC model faithfully reproduces the corresponding trajectories, indicating a reliable long-horizon rollout in stable and periodic regimes. For the chaotic regimes of both maps, we find that the model can still provide a reasonably good forecast for a short horizon, in particular for $k=3$. However, as the horizon is further increased, the predictions become progressively less reliable due to the accumulation of errors and the intrinsic sensitivity of chaotic dynamics.
Figure 9: Closed-loop prediction with different horizons $k$ for logistic map (left) and Henon map (right) for time-steps $t\in[151,200]$. Blue lines represent predicted values and red lines represent actual values. For logistic map $r=2.9$ (fixed point), $r=3.5,3.829$ (periodic orbits) and $r=4$ (chaotic) map parameters have been considered. For Hénon map, a=1.005,1.25a=1.005,1.25(periodic orbits) and $a=1.4$ (chaotic) map parameters have been considered.